OpenStudy (anonymous):

will give medal and become a fan to best answer. solve 4/9 n + 6 = 4/3 A. 10 1/2 B. 16 1/2 C. -4 2/3 D. -10 1/2

OpenStudy (anonymous):

What grade are you in? I might be able to help

OpenStudy (anonymous):

10th @JadeAllen

OpenStudy (anonymous):

Im in the 6th grade sorry.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

a

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

mybad i mean d -10 .5

OpenStudy (anonymous):

umm.... ok....

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

@binks any comment?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

yes d

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(4/9)n +6 -6 = 4/3 -6 (4/9)n=-4 (2/3) (4/9) / (4/9)=-4(2/3) / (4/9) n = -21/2 or -10 1/2